If you invest 35% of your investment in MicroGap with an expected rate of return of 14 % and the remainder in Amgen with an expected rate of return of 8 %, the expected return on your portfolio is:

12.4%

13.6%

10.1%

11.5%
Answer #1

Expected return on portfolio = (weight of investment in MicroGap * expected rate of return in MicroGap) + (weight of investment in Amgen * expected rate of return in Amgen)

weight of investment in Amgen = 1 - 0.35 = 0.65

Expected return on portfolio = (0.35 * 14%) + (0.65 * 8%) = 10.1%